OpenStudy (anonymous):

I have no idea how i am supposed to do this problm. Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.) 2 cos2 θ − sin θ = 1 θ =

OpenStudy (90890908967777878888):

we can factor this: (2sinθ+1)(sinθ−1)=0 sinθ=−12→θ={−π6+2kπ,7π6+2kπ} sinθ=1→θ={π2+2kπ} @Drew1997

OpenStudy (mathmate):

@drew1997 is the question \(2 cos^2(θ) − sin(θ) = 1\)....................(A) or \(2 cos(2θ) − sin(θ) = 1\)....................(B) ?

OpenStudy (anonymous):

That was my mis-type it was supposed to be \[2\cos(\theta)^2-\sin(\theta)=1\]